# LLMs for Generalizable Language-Conditioned Policy Learning under Minimal Data Requirements

### v1

@@ Abstract
To develop autonomous agents capable of executing complex, multi-step decision-making tasks as specified by humans in natural language, existing reinforcement learning approaches typically require expensive labeled datasets or access to real-time experimentation. Moreover, conventional methods often face difficulties in generalizing to unseen goals and states, thereby limiting their practical applicability. This paper presents TEDUO, a novel training pipeline for offline language-conditioned policy learning. TEDUO operates on easy-to-obtain, unlabeled datasets and is suited for the so-called in-the-wild evaluation, wherein the agent encounters previously unseen goals and states. To address the challenges posed by such data and evaluation settings, our method leverages the prior knowledge and instruction-following capabilities of large language models (LLMs) to enhance the fidelity of pre-collected offline data and enable flexible generalization to new goals and states. Empirical results demonstrate that the dual role of LLMs in our framework—as data enhancers and generalizers—facilitates both effective and data-efficient learning of generalizable language-conditioned policies.
@@ 1 Introduction
Motivation. A central aim of AI research is to develop autonomous agents capable of solving complex, multi-step decision-making tasks based on human-provided instructions articulated in natural language. Current approaches, which rely on traditional reinforcement learning (RL) methods, require either vast amounts of data in the form of offline expert demonstrations or data collected from real-time interactions with the environment. Such data is expensive to gather and online experimentation may be impractical in many real-world scenarios. Moreover, even with substantial data, RL agents are often constrained to a limited set of previously attained goals, or their performance deteriorates significantly when tested on new goal-reaching tasks (Yang et al., 2023).
@@
Problem setting. Given the above, this paper approaches the problem of learning generalizable language-conditioned policies under minimal data requirements. Specifically, we consider a fully offline setup with access to a pre-collected dataset of state-action transitions, ${\mathcal{D}}$, alongside an unpaired set of natural language commands, ${\mathcal{G}}^{tr}$. The dataset ${\mathcal{D}}$ consists of triplets $(x,a,x^{\prime})$, where $x$ and $x^{\prime}$ belong to a high-dimensional state space ${\mathcal{X}}$, and $a$ represents actions within an action space ${\mathcal{A}}$. The natural language goals $g\in{\mathcal{G}}^{tr}$ describe a subset of tasks achievable within the environment. We posit that such data is often easy to obtain by simply recording agents interact with the environment and creating a list of natural language commands corresponding to the tasks typically performed within that environment. For instance, in household robotics, ${\mathcal{D}}$ may be derived from random exploration of the environment, while ${\mathcal{G}}^{tr}$ may describe tasks such as “Go and open the window” or “Fetch me a cup of tea.” In the case of personal assistants, ${\mathcal{D}}$ might be collected by recording the daily activities of a human interacting with their mobile device, while ${\mathcal{G}}^{tr}$ would describe goals like “Book a restaurant for 7 PM” or “Send an email to John.” With no assumptions regarding how the offline data has been collected, nor access to the ground-truth state-transition dynamics or rewards, our aim is to learn a language-conditioned policy, $\pi^{*}$, that can determine optimal actions for previously unseen goals $g\notin{\mathcal{G}}^{tr}$ and states $x\notin{\mathcal{D}}$. For a formal definition of the problem setup, please refer to section 2.
@@
Challenges. The task of learning $\pi^{*}$ from ${\mathcal{D}}$ and ${\mathcal{G}}^{tr}$ alone might seem impossible without resorting to human supervision. We can immediately identify the following challenges: C1) Unlabeled data. The dataset ${\mathcal{D}}$ lacks explicit labels linking states $x\in{\mathcal{X}}$ to the goals $g\in{\mathcal{G}}^{tr}$. Nor does it include any rewards indicating the optimality of actions in relation to these goals. C2) Limited exploration. We are in an offline setup with our knowledge of the environment dynamics being constrained to the state-action transitions observed in ${\mathcal{D}}$. C3) Unknown data collection policy. We make no assumptions regarding the optimality of the data collection policy concerning the training or testing goals. The actions in ${\mathcal{D}}$ could be entirely random or generated by policies aimed at solving goals with an unknown relationship to those in ${\mathcal{G}}^{tr}$. C4) Generalization to new goals and states. Beyond solving goals from ${\mathcal{G}}^{tr}$ and taking optimal actions in previously observed states $x\in{\mathcal{D}}$, we want our agent to generalize to new states and language commands corresponding to novel goal states.
@@
Proposed Solution: LLMs to elevate conventional RL. Recent advances in LLMs offer a promising solution to these challenges. LLMs, pre-trained on vast amounts of Internet data, possess the requisite prior knowledge to understand natural language commands and follow simple instructions. However, while LLMs excel at general language comprehension, their ungrounded knowledge is insufficient for executing complex, multi-step decision-making tasks in dynamic environments (Finn, 2024; Szot et al., 2024). In this paper, we propose a novel training pipeline for offline language-conditioned policy learning—TEDUO: Teaching the Environment Dynamics from Unlabeled Observations. As displayed in Figure 1, TEDUO distills knowledge of the environment dynamics into a pre-trained LLM through supervised fine-tuning. This knowledge is obtained by learning optimal policies with traditional RL, based on the offline dataset augmented with LLM-generated state abstractions and labels. Thus, within TEDUO, LLMs fulfill the dual role of cheap data enhancers and flexible generalizers, elevating conventional RL to address challenges C1-C4.
Contributions. 1) We introduce TEDUO—a novel LLM fine-tuning pipeline, which, to the best of our knowledge, is the first to enable the learning of generalizable language-conditioned policies based solely on an unlabeled dataset of state-action transitions and an unpaired set of natural language commands. 2) We demonstrate that LLMs can be effectively employed both as cost-effective data enhancers—automating critical tasks typically handled by humans—and as versatile generalizers—acting as a general language-conditioned policy capable of solving new tasks in previously unseen scenarios. 3) We empirically show that TEDUO enhances both the data efficiency and generalization capacity of offline training, outperforming competing approaches by a significant margin. We analyze the scalability of our method and provide insights into the learning process, showing that fine-tuned LLMs acquire core skills, rather than simply memorize optimal actions.
@@ 2 Problem Formalism
Inputs. We are given a dataset ${\mathcal{D}}$ of past interactions of an agent acting according to a data collection policy $\pi^{\beta}$. This dataset is represented as a collection of trajectories:
${\mathcal{D}}=\{\tau_{i}\}_{i\in{\mathcal{I}}},\tau_{i}=\{(x_{t},a_{t},x_{t+1})\}_{t=0}^{T_{i}},\quad x_{0}\sim\rho,\ x_{t+1}\sim P(\cdot|x_{t},a_{t}),\ a_{t}\sim\pi^{\beta}(\cdot|x_{t}),$ (1)
@@
where $P$ is the state transition function determining the next state given an action $a_{t}\in{\mathcal{A}}$ and state $x_{t}\in{\mathcal{X}}$ and $\rho$ represents a distribution of initial states. Alongside ${\mathcal{D}}$, we are provided with an unpaired set of training goals ${\mathcal{G}}^{tr}$ describing a subset of tasks an agent may attempt to solve within the environment. Each goal $g$ is expressed in a goal representations space ${\mathcal{G}}$. In this paper, we focus on learning language-conditioned policies, taking ${\mathcal{G}}$ to be the space of natural language.
@@
Modeling assumptions. Denoting by $\mathcal{P}({\mathcal{X}})$ the powerset of ${\mathcal{X}}$, we assume there exists a ground-truth mapping $\phi:{\mathcal{G}}\rightarrow\mathcal{P}({\mathcal{X}})$ associating each goal $g$ with a subset of the state space, $\phi(g)={\mathcal{X}}_{g}\subseteq{\mathcal{X}}$. We say that $g$ is achieved at time step $t$, if $x_{t}$ lies in ${\mathcal{X}}_{g}$. [^1] In this paper, we focus on simple goals representable as a subset of the state space. This definition can easily be extended to more complex goals using temporal logic. We leave such extensions for future work.. Then, the cumulative discounted reward: $\sum_{t=0}^{\infty}\gamma^{t}R_{\phi}(x_{t},a_{t},x_{t+1};g)$, with $R_{\phi}(x_{t},a_{t},x_{t+1};g)=\mathbbm{1}\{x_{t+1}\in\phi(g)\}$ measures the optimality of actions taken by an agent with respect to achieving the goal $g$, where $\gamma\in[0,1)$ is the discount factor penalizing long sequences of actions. We make no assumptions regarding the optimality of the data collection policy $\pi^{\beta}$ with respect to ${\mathcal{G}}^{tr}$ and thus, in what follows, we will view our pre-collected data as an un-ordered collection of state-action-state transitions, in short denoted as ${\mathcal{D}}=\{(x,a,x^{\prime})\}$. We also do not assume access to either of the ground-truth state-transition dynamics $P$ or the goal-to-state mapping $\phi$, and consequently the reward $R_{\phi}$. We only require that ${\mathcal{G}}^{tr}$ contains goals corresponding to states that have been visited in ${\mathcal{D}}$  [^2] In practice, ${\mathcal{G}}^{tr}$ can consist of a much larger set of training goals. This set will be effectively reduced to the set of visited goals after the first step of our training pipeline. .
The goal. Given ${\mathcal{D}}$ and ${\mathcal{G}}^{tr}$, our objective is to learn a language-conditioned policy $\pi^{*}$, where
[…]
@@
Crucially, $\pi^{*}$ should generalize to novel goals $g\notin{\mathcal{G}}^{tr}$ and previously unseen states $x\notin{\mathcal{D}}$. We also require that $\pi^{*}$ not only generalizes to synonymous language commands, but also to goals corresponding to new goal-states, emphasizing our focus on evaluation in the wild.
@@ 3 The Method: TEDUO
Step 1. LLMs as data enhancers Construction of abstract MDPs. For each goal, $g\in{\mathcal{G}}^{tr}$, we construct an abstract MDP by employing LLM-automated hindsight labeling and state abstraction, addressing C1 and C2, respectively.
Step 2. Goal-conditioned policy learning with offline RL After obtaining a labeled dataset for each goal in ${\mathcal{G}}^{tr}$, we solve the set of abstract MDPs using an out-of-the-box offline RL algorithm. As a result, we obtain a set of learned policies $\{\pi_{g}:g\in{\mathcal{G}}^{tr}\}$. The learned policies improve on naive imitation learning, addressing C3.
Step 3. LLMs as generalizers LLM supervised fine-tuning. We distill the knowledge about the environment dynamics and optimal actions into a pre-trained LLM with supervised instruction fine-tuning (SFT). This step grounds the prior knowledge of the base LLM in the environment dynamics, thus enabling generalization to new, previously unseen states and goals, addressing both challenges C2 and C4.
In the following paragraphs we explain in detail individual steps of TEDUO.
3.1 Step 1. Construction of abstract MDPs
Starting with the dataset of unlabeled observations ${\mathcal{D}}$ and the training goals ${\mathcal{G}}^{tr}$, we first construct a set of abstract MDPs $\{{\mathcal{M}}^{g}:g\in{\mathcal{G}}^{tr}\}$, where ${\mathcal{M}}^{g}:=({\mathcal{S}}^{g},{\mathcal{A}},P^{g},R^{g},\rho,\gamma)$, with ${\mathcal{S}}^{g}$ being the abstract state space for the goal $g$, $P^{g}$ the induced transition operator, and $R^{g}$ the reward function with respect to the goal $g$. To define our abstract MDPs, we employ hindsight labeling and state abstraction with LLM-based operators.
3.1.1 State abstraction
The goal of state abstraction is to reduce the size of the environmental state space by grouping together similar states in a way that reduces the complexity of the underlying problem being solved (Li et al., 2006). With a well-designed state abstraction, RL algorithms can learn more efficiently, requiring fewer samples of data, which is particularly relevant in our offline setup. Formally, let $F:{\mathcal{X}}\times{\mathcal{G}}\rightarrow{\mathcal{S}}^{g}$ be an abstraction operator that given a goal $g$ maps a single observation $x\in{\mathcal{X}}$ to an abstract state $s^{g}=F(x;g)$. The abstraction map, $F$, should be such that $|{\mathcal{X}}|\gg|{\mathcal{S}}^{g}|$.
@@ 3.1.1 State abstraction
In this paper, we use natural language to guide state abstraction, so that the resulting abstract states contain only the goal-relevant information. Namely, we consider environments that can be represented in a $d$-dimensional feature space, $\mathcal{X}=\mathcal{X}^{1}\times\mathcal{X}^{2}\times\ldots\times\mathcal{X}^{d}$. We assume that only a relatively small subset of these variables is relevant for solving a specific goal $g$ and an even smaller subset is required to identify the goal states $\phi(g)$.
The state abstraction operator is implemented as a collection of Python functions built on top of the feature selection made by a prompted language model, $F(\ \cdot\ ;g)=\mathrm{LLM}^{abstrct}(g)(\cdot)$. Using LLM powered Python functions instead of directly applying the LLM to create an abstraction of each state reduces the number of LLM calls from $|\mathcal{X}||G^{tr}|$ to $|G^{tr}|$ and ensures that the abstraction is consistent across all states. The prompt for generating the code includes contextual information about the environment, a list of features, and a description of the goal, instructing the LLM to create a function that removes features of a state that are irrelevant to achieving the specified goal. Additionally, state abstraction functions separate the set of relevant features into two subsets: $s^{g}_{\phi}$ and $s^{g}_{\bar{\phi}}$, so that $s^{g}_{\phi}\cup s^{g}_{\bar{\phi}}$. The features in $s^{g}_{\phi}$ are the ones which are necessary for identifying if the underlying low-level state achieves $g$, i.e. if $x\in\phi(g)$. The remaining features, which are relevant for solving the task specified by $g$ but not strictly necessary for identifying if this goal is achieved are found in $s^{g}_{\bar{\phi}}$. We introduce this separation of abstracted features to even further reduce the dimensionality of the state space for hindsight labeling (see next section). Figure 2 shows an example effect of applying our LLM state abstraction on a grid-world from the BabyAI environment.
@@
Following recent works (Kwon et al., 2023), we hypothesize that the existing abilities of LLMs in natural language understanding are sufficient to perform the simple task of identifying whether a particular state belongs to the set of goal states $\phi(g)$ associated with a given goal description $g$. In order to perform hindsight labeling of our dataset ${\mathcal{D}}$, we wish to approximate $\phi$, and thus the reward $R_{\phi}(\cdot;g)$, with a prompted language model $\mathrm{LLM}^{rwrd}\approx\mathbbm{1}_{(\text{$g$ is achieved in $s^{g}$})}$. Note, we assign the rewards in the abstracted spaces ${\mathcal{S}}^{g}$ and not the original state space ${\mathcal{X}}$. Given the large number of goals and states, to reduce the number of LLM calls needed, instead of using language models directly, we train proxy reward models–lightweight neural networks trained to predict the labels generated by the prompted language model, $\mathrm{LLM}^{rwrd}$. For each goal $g$, we build a supervised dataset, $\{(s^{g},r^{g}):r^{g}=\mathrm{LLM}^{rwrd}(s^{g};g),s^{g}\in\tilde{{\mathcal{S}}}^{g}\}$, where $\tilde{{\mathcal{S}}}^{g}$ is a small, diverse subset of the abstract space ${\mathcal{S}}^{g}$ and $r^{g}\in\{0,1\}$. We train a neural network $R_{\theta}(\ \cdot\ ;g):{\mathcal{S}}^{g}\rightarrow\{0,1\}$ to predict binary rewards for the entire abstract state space ${\mathcal{S}}^{g}$. The subset $\tilde{{\mathcal{S}}}^{g}$ is chosen so that for any two abstract states, the set of features relevant for goal-identifications is distinct, i.e. $\forall s^{g}_{1}\neq s^{g}_{2}\in\tilde{{\mathcal{S}}}^{g},s^{g}_{1,\phi}\neq s^{g}_{2,\phi}$. This maximizes the chances of including goal-states in $\tilde{{\mathcal{S}}}^{g}$, mitigating the potential issue of generating a highly-imbalanced dataset for training our proxy neural networks. These proxy reward functions provide a much more cost-effective way to perform hindsight labeling compared to labeling all states from ${\mathcal{D}}$ for all goals from ${\mathcal{G}}^{tr}$ directly by LLM prompting or with human annotators. Appendix D shows that for the BabyAI environment, depending on the goal, proxy rewards reach near 100% accuracy in comparison to the ground truth rewards of the environment.
3.2 Step 2. Solving the Abstract MDPs
@@
After applying state abstractions and hindsight labeling to our offline dataset ${\mathcal{D}}$, for each goal $g\in{\mathcal{G}}^{tr}$, we obtain an offline dataset ${\mathcal{D}}^{g}:=\{(s^{g},a,s^{g},r^{g})\}$. Given these data, we can apply any offline reinforcement learning method to learn optimal policies $\pi^{g}$, for each goal $g\in{\mathcal{G}}^{tr}$. In practice, however, to learn the goal-conditioned policies, the chosen RL method should be scalable, as we must solve multiple MDPs, one for each goal in ${\mathcal{G}}^{tr}$. Therefore, in our instantiation, we discard computationally intensive methods. Furthermore, as the generalization to new states is tackled by the next step, we do not require at this stage that the learned policies generalize to unseen states. Given these considerations, we simply choose tabular Q-learning (Watkins & Dayan, 1992) to solve the set of abstract MDPs. At the end of this stage, we obtain a set of learned policies $\{\pi^{g}:g\in{\mathcal{G}}^{tr}\}$. These policies are limited to the set of training goals and the set of states observed in ${\mathcal{D}}$. The final step of our pipeline addresses these limitations.
3.3 Step 3. Training the LLM as a goal-conditioned policy
@@
To enable generalization to previously unseen states, and more importantly, generalization to novel goals, the final step of our method distills the knowledge of the optimal actions per each abstract state and goal into a pre-trained LLM. We build a supervised dataset ${\mathcal{D}}^{SFT}$ consisting of goal commands, initial abstract states and the sequence of optimal actions with respect to the learned policies. Concretely, we have
${\mathcal{D}}^{SFT}:=\{(g,s_{0}^{g},[a_{0}^{*,g},\ldots,a_{n_{g}}^{*,g}])$ $:\quad g\in{\mathcal{G}}^{tr},\ s_{0}^{g}\in{\mathcal{D}}^{g},\ a_{t}^{*,g}=\operatorname*{arg\,max}_{a\in{\mathcal{A}}}\pi^{g}(a\mid s_{t}^{g}),$
$s_{t+1}^{g}=\operatorname*{arg\,max}_{s\in{\mathcal{S}}^{g}}\hat{P}^{g}(s|s_{t}^{g},a_{t}^{*,g})\},\ n_{g}\ s.t.\ R_{\hat{\theta}}(s_{n_{g}+1}^{g};g)=1\},$

### v2

# The Synergy of LLMs & RL Unlocks Offline Learning of Generalizable Language-Conditioned Policies with Low-fidelity Data

@@ Abstract
Developing autonomous agents capable of performing complex, multi-step decision-making tasks specified in natural language remains a significant challenge, particularly in realistic settings where labeled data is scarce and real-time experimentation is impractical. Existing reinforcement learning (RL) approaches often struggle to generalize to unseen goals and states, limiting their applicability. In this paper, we introduce TEDUO, a novel training pipeline for offline language-conditioned policy learning in symbolic environments. Unlike conventional methods, TEDUO operates on readily available, unlabeled datasets and addresses the challenge of generalization to previously unseen goals and states. Our approach harnesses large language models (LLMs) in a dual capacity: first, as automatization tools augmenting offline datasets with richer annotations, and second, as generalizable instruction-following agents. Empirical results demonstrate that TEDUO achieves data-efficient learning of robust language-conditioned policies, accomplishing tasks beyond the reach of conventional RL frameworks or out-of-the-box LLMs alone.
Machine Learning, ICML
@@ 1 Introduction
Motivation. Enabling AI agents to follow human-provided natural language instructions is a critical step toward building intelligent systems that can perform complex, multi-step tasks in dynamic environments. However, existing methods often require vast quantities of annotated data or rely on costly online interactions with the environment to train effective policies. Such requirements limit their scalability to real-world scenarios where collecting annotated or interaction data is infeasible. Moreover, current RL-based agents often struggle to generalize to new goals and environments beyond their training data (Yang et al., 2023). These challenges underline the need for a change of focus in the area of instruction-following agents: from reliance on online interactions in environments with rich forms of feedback to the learning of generalizable policies from unlabeled observational data. We illustrate the problem setting considered with the following hypothetical scenario:
Imagine recording a human using their smartphone in daily tasks, such as text messaging or calendar management. How can we train an an AI assistant following natural language instructions, e.g.“Book a restaurant for 7 PM” or “Send an email to John.”, by just recording the actions taken by the human and not having to manually label each state or action recorded? How can we make the learned policies generalize to commands that have not been explicitly observed in the data?
@@
Problem setting. We formalize our setup by framing it as an offline language-conditioned policy learning problem. We assume access to a pre-collected dataset of state-action transitions, ${\mathcal{D}}$ consisting of triplets $(x,a,x^{\prime})$, where $x$ and $x^{\prime}$ belong to an observable state space ${\mathcal{X}}$, and $a$ represents actions within an action space ${\mathcal{A}}$. To enable the grounding of the environment dynamics with the space of natural language we additionally require: 1) that the individual states and actions are representable in a textual format; 2) access to an unordered list of goals, ${\mathcal{G}}$ expressed in natural language that are plausibly achievable within the environment of interest. We posit that both ${\mathcal{D}}$ and ${\mathcal{G}}$ are often easy to obtain by simply recording agents interact with the environment and curating a list of natural language commands corresponding to the tasks typically performed within that environment. Without constraining assumptions regarding the optimality of the data collection policy with respect to the goals ${\mathcal{G}}$, nor access to the ground-truth state-transition dynamics or rewards, our aim is to learn a language-conditioned policy, $\pi^{*}$, that can determine optimal actions with respect to new states $x\notin{\mathcal{D}}$ and new goals $g^{*}\not\in{\mathcal{G}}^{tr}$, where ${\mathcal{G}}^{tr}\subset{\mathcal{G}}$ is the subset of goals used for training (see section 2 for details).
@@
Challenges. The task of learning $\pi^{*}$ from ${\mathcal{D}}$ and ${\mathcal{G}}^{tr}$ alone might seem impossible without resorting to human supervision. We can immediately identify the following challenges: C1) Unlabeled data. The dataset ${\mathcal{D}}$ lacks explicit labels linking states $x\in{\mathcal{X}}$ to the goals $g\in{\mathcal{G}}$. Nor does it include any rewards indicating the optimality of actions in relation to these goals. C2) Limited exploration. We are in an offline setup with our knowledge of the environment dynamics being constrained to the state-action transitions observed in ${\mathcal{D}}$. C3) Unknown data collection. We make no assumptions regarding the optimality of the data collection policy with respect to the training or testing goals. The actions in ${\mathcal{D}}$ could be entirely random or generated by policies aimed at solving goals with an unknown relationship to those in ${\mathcal{G}}$. C4) Generalization. Beyond solving goals from ${\mathcal{G}}^{tr}$ and taking optimal actions in previously observed states $x\in{\mathcal{D}}$, we want our agent to generalize to new states and language commands corresponding to genuinely novel goal states.
@@
Solution: LLMs & RL synergy. Recent advances in LLMs offer a promising solution to these challenges. LLMs, pre-trained on vast amounts of Internet data, possess the requisite prior knowledge to understand natural language and follow simple instructions. However, while LLMs excel at general language comprehension, their ungrounded knowledge is insufficient for executing complex, multi-step decision-making tasks in dynamic environments (Finn, 2024; Szot et al., 2024). In this paper, we propose a novel, sequential training pipeline for offline language-conditioned policy learning—TEDUO: Teaching the Environment Dynamics from Unlabeled Observations. As displayed in Figure 1, TEDUO distills knowledge of the environment dynamics into a pre-trained LLM through supervised fine-tuning (SFT). This knowledge is obtained by learning optimal policies with traditional RL, based on the offline dataset augmented with LLM-generated labels and optional state abstractions. Within TEDUO, LLMs fulfill the dual role of cheap data enhancers and flexible generalizers, elevating conventional RL to address challenges C1-C4.
Contributions. We categorize our contributions as having a significant impact both for the fields of LLMs and RL.
[…]
@@ 2 Problem Formalism
We are given a dataset ${\mathcal{D}}$ of past interactions of an agent acting according to a data collection policy $\pi^{\beta}$. This dataset is represented as a collection of trajectories:
${\mathcal{D}}=\{\tau_{i}\}_{i\in{\mathcal{I}}},\ \tau_{i}=\{(x_{t},a_{t},x_{t+1})\}_{t=0}^{T_{i}},$
$x_{0}\sim\rho,\ x_{t+1}\sim P(\cdot|x_{t},a_{t}),\ a_{t}\sim\pi^{\beta}(\cdot|x_{t}),$
@@
where $P$ is the state transition function determining the next state given an action $a_{t}\in{\mathcal{A}}$ and state $x_{t}\in{\mathcal{X}}$ and $\rho$ represents the distribution of initial states. Alongside ${\mathcal{D}}$, we are given an unpaired set of goals ${\mathcal{G}}$ split into training ${\mathcal{G}}^{tr}$ and testing ${\mathcal{G}}^{test}$ subsets, describing in natural language tasks an agent may attempt to solve within the environment.
@@
Denoting by $\mathcal{P}({\mathcal{X}})$ the powerset of ${\mathcal{X}}$, we assume there exists a ground-truth mapping $\phi:{\mathcal{G}}\rightarrow\mathcal{P}({\mathcal{X}})$ associating each goal $g$ with a subset of the state space, $\phi(g)={\mathcal{X}}_{g}\subseteq{\mathcal{X}}$. We say that $g$ is achieved at time step $t$, if $x_{t}$ lies in ${\mathcal{X}}_{g}$. [^1] In this paper, we focus on simple goals representable as a subset of the state space. This can be extended to more complex goals using temporal logic which we leave for future work.. Then, the cumulative discounted reward: $\sum_{t=0}^{\infty}\gamma^{t}R_{\phi}(x_{t},a_{t},x_{t+1};g)$, with $R_{\phi}(x_{t},a_{t},x_{t+1};g)=\mathbbm{1}\{x_{t+1}\in\phi(g)\}$ measures the optimality of actions taken by an agent with respect to achieving the goal $g$, where $\gamma\in[0,1)$ is the discount factor penalizing long sequences of actions. We make no assumptions regarding the optimality of the data collection policy $\pi^{\beta}$ with respect to ${\mathcal{G}}^{tr}$ and thus, in what follows, we will view our pre-collected data as an un-ordered collection of state-action-state transitions, in short denoted as ${\mathcal{D}}=\{(x,a,x^{\prime})\}$. We also do not assume access to either of the ground-truth state-transition dynamics $P$ or the goal-to-state mapping $\phi$, and consequently the reward $R_{\phi}$. We only require that ${\mathcal{G}}^{tr}$ contains goals corresponding to states that have been visited in ${\mathcal{D}}$ [^2] In practice, ${\mathcal{G}}^{tr}$ can consist of a much larger set of training goals. This set will be effectively reduced to the set of visited goals after the first step of our training pipeline..
@@
The goal. Given ${\mathcal{D}}$ and ${\mathcal{G}}^{tr}$, our objective is to learn a language-conditioned policy $\pi^{*}$ maximizing the expectation of the cumulative discounted rewards averaged across all $g\in{\mathcal{G}}$. Crucially, $\pi^{*}$ should generalize to novel goals $g\notin{\mathcal{G}}^{tr}$ and previously unseen states $x\notin{\mathcal{D}}$. We require that $\pi^{*}$ not only generalizes to synonymous language commands, but also to previously unvisited goal-states.
@@ 3 The Method: TEDUO
[…]
    Construction of solvable MDPs. For each goal, $g\in{\mathcal{G}}^{tr}$, we construct an MDP by employing LLM-automated hindsight labeling and optional state abstraction, addressing C1 and C2.
[…]
    Offline Policy Learning. After obtaining a labeled dataset for each goal in $g\in{\mathcal{G}}^{tr}$, we solve the set of abstract MDPs using an out-of-the-box offline RL algorithm. As a result, we obtain a set of learned policies $\{\pi_{g}:g\in{\mathcal{G}}^{tr}\}$. The learned policies improve on naive imitation learning, addressing C3.
[…]
    LLM supervised fine-tuning. We distill the knowledge of the environment dynamics into a pre-trained LLM via SFT, grounding the prior knowledge of the base LLM and thus, enabling generalization to new, previously unseen states and goals, addressing C2 and C4.
3.1 Step 1. Construction of solvable MDPs
Given that in many symbolic environments the original state representations ${\mathcal{X}}$ is high-dimensional, our pipeline contains an optional state abstraction step that maps raw observations $x\in{\mathcal{X}}$ to a goal-dependent abstract states $s^{g}\in{\mathcal{S}}^{g}$. For each $g\in G^{tr}$ we construct an MDP $\{{\mathcal{M}}^{g}:g\in{\mathcal{G}}^{tr}\}$, where ${\mathcal{M}}^{g}:=({\mathcal{S}}^{g},{\mathcal{A}},P^{g},R^{g},\rho,\gamma)$, with $P^{g}$ standing for the induced transition operator and $R^{g}$ the reward function with respect to the goal $g$ approximated with scalable LLM-based proxies. The state abstraction stage is entirely optional and if not employed, we simply take ${\mathcal{S}}^{g}={\mathcal{X}}$ and $P^{g}=P$.
3.1.1 State abstraction (optional)
The goal of state abstraction is to reduce the size of the state space by grouping together similar states in a way that reduces the complexity of the MDP (Li et al., 2006). With a well-designed abstraction, conventional RL algorithms can learn more efficiently, requiring fewer samples of data, which is particularly relevant in the offline setup. Let $F:{\mathcal{X}}\times{\mathcal{G}}\rightarrow{\mathcal{S}}^{g}$ be an abstraction operator that given a goal $g$ maps a single observation $x\in{\mathcal{X}}$ to an abstract state $s^{g}:=F(x;g)$. This map should be such that $|{\mathcal{S}}^{g}|\ll|{\mathcal{X}}|$.
@@ 3.1.1 State abstraction (optional)
In the context of text-based environments, we can leverage natural language to guide state abstraction, so that the resulting abstract states contain only the goal-relevant information. Given an environment that can be essentially represented in a $d$-dimensional feature space, $\mathcal{X}=\mathcal{X}^{1}\times\mathcal{X}^{2}\times\ldots\times\mathcal{X}^{d}$, we assume that only a relatively small subset of these variables is relevant for solving a specific goal $g$ and an even smaller subset is required to identify the goal states $\phi(g)$. For details on the implementation of the LLM state abstraction operator refer to Appendix C.4. Figure 2 shows an example effect of state abstraction on a grid-world from the BabyAI environment.
@@
Following recent works (Kwon et al., 2023), we hypothesize that the existing abilities of LLMs are sufficient to perform the simple task of identifying whether a particular state belongs to the set of goal states $\phi(g)$ associated with the goal description $g$. In order to perform hindsight labeling of our dataset ${\mathcal{D}}$, we wish to approximate $\phi$, and thus the reward $R_{\phi}(\cdot;g)$, with a prompted language model $\mathrm{LLM}^{rwrd}\approx\mathbbm{1}_{(\text{$g$ is achieved in $s^{g}$})}$, where we assign the rewards in the abstracted spaces ${\mathcal{S}}^{g}$ instead of ${\mathcal{X}}$. Given the large number of goals and states, to reduce the number of LLM calls needed, instead of using language models directly, we train proxy reward models–lightweight neural networks $R_{\theta}(\ \cdot\ ;g):{\mathcal{S}}^{g}\rightarrow\{0,1\}$–to predict the labels generated by the prompted language model, $\mathrm{LLM}^{rwrd}$. Details can be found in the Appendix C.5. The proxy rewards functions provide a much more cost-effective way to perform hindsight labeling compared to labeling all states from ${\mathcal{D}}$ for all goals from ${\mathcal{G}}^{tr}$ directly by LLM prompting let alone by human annotators. Appendix B.5 shows that for the BabyAI environment proxy rewards reach near 100% accuracy in comparison to the ground truth rewards of the environment.
3.2 Step 2. Offline Policy Learning.
@@
After the first step, for each goal $g\in{\mathcal{G}}^{tr}$ we obtain an offline dataset ${\mathcal{D}}^{g}:=\{(s^{g},a,s^{g},r^{g})\}$. Given these data, we can apply any offline reinforcement learning method to learn optimal policies $\pi^{g}$, for each goal $g\in{\mathcal{G}}^{tr}$. In practice, however, to learn the goal-conditioned policies, the chosen RL method should be scalable, as we must solve multiple MDPs, one for each goal in ${\mathcal{G}}^{tr}$. Therefore, in our instantiation, we discard computationally intensive methods. Furthermore, as the generalization to new states is tackled by the next step, we do not require at this stage that the learned policies generalize to unseen states. Given these considerations, we primarily employ tabular Q-learning (Watkins & Dayan, 1992) to solve the set of abstract MDPs. To demonstrate the flexibility of our approach, we also conduct experiments with Deep Q-Learning (Appendix B.4) and filtered Behavioral Cloning (Appendix B.1). At the end of this stage, we obtain a set of learned policies $\{\pi^{g}:g\in{\mathcal{G}}^{tr}\}$. These policies are limited to the set of training goals and the set of states observed in ${\mathcal{D}}$.
3.3 Step 3. Training the LLM as a generalizable policy
@@
To enable generalization to previously unseen states, and more importantly, generalization to novel goals, the final step of our method distills the knowledge of the optimal actions per each abstract state and goal into a pre-trained LLM. We build a supervised dataset ${\mathcal{D}}^{SFT}$ consisting of goal commands, initial abstract states and the sequence of optimal actions with respect to the learned policies. Concretely,
${\mathcal{D}}^{SFT}:=$$\{(g,s_{0}^{g},[a_{0}^{*,g},\ldots,a_{n_{g}}^{*,g}]):g\in{\mathcal{G}}^{tr},\ s_{0}^{g}\in{\mathcal{D}}^{g},$
$a_{t}^{*,g}=\operatorname*{arg\,max}_{a\in{\mathcal{A}}}\pi^{g}(a\mid s_{t}^{g}),$
$s_{t+1}^{g}=\operatorname*{arg\,max}_{s\in{\mathcal{S}}^{g}}\hat{P}^{g}(s|s_{t}^{g},a_{t}^{*,g})\},$
$n_{g}\ s.t.\ R_{\hat{\theta}}(s_{n_{g}+1}^{g};g)=1\},$